40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Dec 11 09:31:47 2012, Ayaka, Update, LSC, Oplev update for improving sensitivity 7x
    Reply  Tue Dec 11 21:53:37 2012, Ayaka, Update, LSC, Oplev update for improving sensitivity OL_BSpit.pdfBSoplevservo.png
Message ID: 7808     Entry time: Tue Dec 11 09:31:47 2012     Reply to this: 7812
Author: Ayaka 
Type: Update 
Category: LSC 
Subject: Oplev update for improving sensitivity 


We observed that oplev servos affect the arm spectra badly (elog #7798). Some of them are fixed, but still they inject noise into the arms.
So I tried to turn the oplevs off and to see the acoustic noise effect. However, the mirrors moves so much that the signal does not seem to be linear any more, and the noise spectrum of arms changes especially around 60 - 100 Hz as you can see the spectrogram of YARM error signal below. This makes it difficult to find acoustic coupling noise. Therefore, I tried to fix the oplev servos so that the noise spectra do not get worse when the oplev servos are on.

Checking oplev UGFs

I checked the oplev open loop transfer functions. The UGFs of oplevs are all around 1-3Hz and phase margin looks enough except the BS oplev.
The gain of the BS oplev OLTF has so low that the signal is not fed back. Moreover, there is much phase delay in the BS feedback loop than the others'.
The counts of BS oplev sum is not changed so much for this 4 months, so the oplev beam seems to hit the BS correctly.
I am not sure what makes difference.



Clipped oplev beam fixed

Den and I found the output beam of ETMY oplev was clipped the other day. Also I found the scattered beam of ITMY oplev was on the edge of the mirror inside the vacuum and it made more scattered lights.
ITMYoplev_before.jpg(before) -> IMG_0128.JPG(after)

  I fixed both of the clipped beam. But still the oplev feedback inject the noise into the arm. (red: oplev off, blue: oplev on)

ELOG V3.1.3-